If 3 points of yI are 13,19,4 and y2 are17,9,10 AND y3 are 128 ,32 find 3rd

Grade:12th pass

2 Answers

Arun
3 years ago
Dear Manish
 
Please elaborate the question. if possible, pleattach an image, I will be happy to help you.
 
Regards
Vikas TU
3 years ago
Dear student 
Given  
if 3 points of y I are 13,19,4 and y 2  are 17,9,10 AND y 3 are 128 ,32 find 3rd
.The given series is  
13    19    4
 71     9     10
 128    32    ?
So it can be written as  
13 + 19 = 32 / 4 = 8
71 + 9 = 80 / 10 = 8
128 + 32 = 160 / 20 = 8
So the missing number will be 20